# The Equational Theories Project: Advancing Collaborative Mathematical Research at Scale

### v1

@@ 1. Introduction
The purpose of this paper is to report on the Equational Theories Project (ETP) [^1] https://teorth.github.io/equational_theories/, a pilot project launched [^2] https://terrytao.wordpress.com/2024/09/25 in September 2024 to explore new ways to collaboratively work on mathematical research projects using machine assistance. The project goal, in the area of universal algebra, was selected [^3] The specific mathematical goal was inspired by the MathOverflow question “Is there an identity between the associative identity and the constant identity?”, posed on July 17, 2023. to be particularly amenable to crowdsourced and computer-assisted techniques, while still being of mathematical research interest.
@@ 1.1. Magmas and equational laws
An equational theory is a collection of equational laws; in this paper we will primarily be concerned with theories generated by a single such law, although it is certainly of interest to explore larger theories as well. Equational theories are one of the simplest non-trivial examples of a theory in the model-theoretic sense; they also arise in various areas of computer science, such as term rewriting systems [term-rewriting], automated theorem proving [mccune-survey], and in the Dolev–Yao model [dolev] of interactive cryptographic protocols.
@@
We also define a variant: we say that $\mathrm{E}$ entails $\mathrm{E}^{\prime}$} for finite magmas, and write ${\mathrm{E}}\models_{\mathrm{fin}}{\mathrm{E}^{\prime}}$, if every finite magma that satisfies $\mathrm{E}$, also satisfies $\mathrm{E}^{\prime}$}. Clearly, the relation ${\mathrm{E}}\models{\mathrm{E}^{\prime}}$ implies ${\mathrm{E}}\models_{\mathrm{fin}}{\mathrm{E}^{\prime}}$; but, as observed by Austin [austin_finite], the converse is not true in general.
@@
  • The central groupoid law $\mathrm{x}\simeq(\mathrm{y}\diamond\mathrm{x})\diamond(\mathrm{x}\diamond\mathrm{z})$ (E168) is an order-$3$ law introduced by Evans [evans] and studied further by Knuth [knuth] and many further authors, being closely related to central digraphs (also known as unique path property digraphs), and leading in particular to the discovery of the Knuth-Bendix algorithm [knuth-bendix]; see [klt] for a more recent survey.
@@
  • Tarski’s axiom $\mathrm{x}\simeq\mathrm{y}\diamond(\mathrm{z}\diamond(\mathrm{x}\diamond(\mathrm{y}\diamond\mathrm{z})))$ (E543) is an order-$4$ law that was shown by Tarski [Tarski1938] to characterize the operation of subtraction in an abelian group; that is to say, a magma ${\mathcal{M}}=(M,\diamond)$ satisfies $\mathrm{E}543$ if and only if there is an abelian group structure on $\mathcal{M}$} for which $x\diamond y=x-y$ for all $x,y\in M$.
@@
  • In a similar vein, it was shown in [mendelsohn-padmanabhan] (see also [meredith-prior]) that the order-$4$ law $\mathrm{x}\simeq(\mathrm{y}\diamond\mathrm{z})\diamond(\mathrm{y}\diamond(\mathrm{x}\diamond\mathrm{z}))$ (E1571) characterizes addition (or subtraction) in an abelian group of exponent $2$; it was shown in [mccune_et_al] that the order-$6$ law $\mathrm{x}\simeq(\mathrm{y}\diamond((\mathrm{x}\diamond\mathrm{y})\diamond\mathrm{y}))\diamond(\mathrm{x}\diamond(\mathrm{z}\diamond\mathrm{y}))$ (E345169) characterizes the Sheffer stroke in a boolean algebra, and it was shown in [higman-neumann] that the order-$8$ law $\mathrm{x}\simeq\mathrm{y}\diamond((((\mathrm{y}\diamond\mathrm{y})\diamond\mathrm{x})\diamond\mathrm{z})\diamond(((\mathrm{y}\diamond\mathrm{y})\diamond\mathrm{y})\diamond\mathrm{z}))$ (E42323216) characterizes division in a (not necessarily abelian) group.
@@
Some further examples of laws characterizing well-known algebraic structures are listed in [mccune-survey].
@@
The Birkhoff completeness theorem [term-rewriting, Th. 3.5.14] implies that an implication ${\mathrm{E}}\models{\mathrm{E}^{\prime}}$ of equational laws holds if and only if the left-hand side of $\mathrm{E}^{\prime}$} can be transformed into the right-hand side by a finite number of substitution rewrites using the law $\mathrm{E}$. However, the problem of determining whether such an implication holds is undecidable in general [mckenzie]. Even when the order is small, some implications [^4] Another contemporaneous example of this phenomenon was the solution of the Robbins problem [robbins]. can require lengthy computer-assisted proofs; for instance, it was noted in [Kisielewicz2] that the order-$4$ law $\mathrm{x}\simeq(\mathrm{y}\diamond\mathrm{x})\diamond((\mathrm{x}\diamond\mathrm{z})\diamond\mathrm{z})$ (E1689) was equivalent to the singleton law $\mathrm{x}\simeq\mathrm{y}$ (E2), but all known proofs were found with computer assistance. [^5] We improved such a proof to make it human-readable, see the blueprint of the ETP. Furthermore, for the finite magma implication relation ${\mathrm{E}}\models_{\mathrm{fin}}{\mathrm{E}^{\prime}}$, no analogue of the Birkhoff completeness theorem is available.
@@ 1.2. The Equational Theories Project
Such systematic determinations of implication graphs have been seen previously in the literature; for instance, in [phillips-vojtechovsky], the relations between $60$ identities of Bol–Moufang type were established, and in the blog post [Wolfram_2022, §17], some initial steps towards generating this graph for the first hundred or so laws on our list were performed. However, to our knowledge, the ETP is the first project to study such implications at the scale of thousands of laws.
The ETP requires the determination of the truth or falsity of $4694$^{2}=$22\,033\,636$$ implications (for both arbitrary magmas and finite magmas), or $4694$\times($4694$-1)=$22\,028\,942$$ if the reflexive implications ${\mathrm{E}}\models\mathrm{E}$ are removed; while one can use properties such as the transitivity of entailment to reduce the work somewhat, this is clearly a task that requires significant automation. It was also a project highly amenable to crowdsourcing, in which different participants could work on developing different techniques, each of which could be used to fill out a different part of the implication graph. In this respect, the project could be compared with a Polymath project [Gowers2009], which used online forums such as blogs and wikis to openly collaborate on a mathematical research problem. However, the Polymath model required human moderators to review and integrate the contributions of the participants, which clearly would not scale to the ETP which required the verification of over twenty million mathematical statements. Instead, the ETP was centered around a GitHub repository in which the formal mathematical contributions had to be entered in the proof assistant language Lean, where they could be automatically verified. In this respect, the ETP was more similar to the recently concluded Busy Beaver Challenge [^6] https://bbchallenge.org/, which was a similarly crowdsourced project that computed the fifth Busy Beaver number $BB(5)$ to be 47 176 87047\,176\,870 through an analysis of about $180$ million Turing machines, with the halting analysis being verified in a variety of computer languages, with the final formal proof written in the proof assistant language Coq [the_coq_development_team_2024_14542673, bbchallenge_bb5]. One of the aims of the ETP was to explore potential workflows for such collaborative, formally verified mathematical research projects that could serve as a model for future projects of this nature.
@@ 1.3. Outcomes
In the course of completing the implication graph, some interesting new algebraic structures were discovered. One such example concerns the magmas satisfying $\mathrm{E}1485$, which we refer to as weak central groupoids as they contain the central groupoids (satisfying $\mathrm{E}168$) as a subclass. In [knuth] it was observed that all finite central groupoids have order equal to a perfect square $n^{2}$; empirically, we have found that finite weak central groupoids always have order $n^{2}$ or $2n^{2}$, although we have no rigorous proof of this claim; they also have a graph-theoretic interpretation analogous to the interpretation of central groupoids as digraphs with the unique path property. For these and other observations we refer the reader to the blueprint of the ETP.
@@ 2. Notation and mathematical foundations
A theory is a collection $\Gamma$ of equational laws; we say that a magma $\mathcal{M}$} satisfies a theory, and write ${\mathcal{M}}\models\Gamma$, if every law in $\Gamma$ is satisfied by $\mathcal{M}$}. If $\mathrm{E}$ is an equational law, we write $\Gamma\models\mathrm{E}$ if every magma that satisfies $\Gamma$ also satisfies $\mathrm{E}$. A free magma ${\mathcal{M}}_{X,\Gamma}=(M_{X,\Gamma},\diamond)$ for such a theory $\Gamma$ and an alphabet $X$ is a magma satisfying $\Gamma$ together with a map $\iota_{X,\Gamma}\colon X\to M_{X,\Gamma}$ which is universal in the sense that every function $f\colon X\to{\mathcal{M}}$ to a magma $\mathcal{M}$} satisfying $\Gamma$ uniquely determines a homomorphism $\varphi_{f,\Gamma}\colon{\mathcal{M}}_{X,\Gamma}\to{\mathcal{M}}$ such that $\varphi_{f,\Gamma}\circ\iota_{X,\Gamma}=f$. This magma is unique up to isomorphism; a canonical way to construct it is as the quotient ${\mathcal{M}}_{X}/\sim_{\Gamma}$ of the free magma ${\mathcal{M}}_{X}$ by the equivalence relation $\sim_{\Gamma}$ given by declaring $w\sim_{\Gamma}w^{\prime}$ if $\Gamma\models w\simeq w^{\prime}$ [term-rewriting, Theorem 3.5.6]. If $\Gamma=\{\mathrm{E}\}$ consists of a single law $\mathrm{E}$, we write ${\mathcal{M}}_{X,\mathrm{E}}$, $\sim_{\mathrm{E}}$, $\varphi_{f,\mathrm{E}}$ for ${\mathcal{M}}_{X,\{\mathrm{E}\}}$, $\sim_{\{\mathrm{E}\}}$, $\varphi_{f,\{\mathrm{E}\}}$ respectively.
@@ 3. Formal foundations
All proofs in the ETP were ultimately formalized in the proof assistant language Lean, though in many cases the proofs were first written in an informal human document, which was then incorporated into the human-readable blueprint [leanblueprint] that accompanied the formalization. Many of the computer-assisted proofs were also first generated as computer output from a source other than Lean, such as an ATP, and later converted to a Lean proof by a separate program custom-written for this task.
@@
@[equational_result]theorem _root_.Equation1437_not_implies_Equation4269 $\exists$ (G : Type) (_ : Magma G), Equation1437 G $\wedge$ $\neg$ Equation4269 G := by use $\mathbb{N}$ × Fin 3, ⟨op⟩ constructor · intro x y z simp [op, add_assoc] · simp only [not_forall, op] use (0, 0), (2, 0) decide
@@
@[equational_result]theorem «Facts from All4x4Tables [[1,2,3,4,5,0],[4,1,2,5,0,3],[3,0,5,2,1,4],[0,5,4,3,2,1],[5,4,1,0,3,2],[2,3,0,1,4,5]]» : $\exists$ (G : Type) (_ : Magma G) (_: Finite G), Facts G [1316, 2863] [411, 680, 817, 1020, 1426, 2035, 2441, 2644, 2853, 2855, 2865, 2872, 2947, 3050, 3253, 3456, 4270, 4283, 4290, 4380, 4598, 4605, 4656] :=⟨Fin 6, «All4x4Tables [[1,2,3,4,5,0],[4,1,2,5,0,3],[3,0,5,2,1,4],[0,5,4,3,2,1],[5,4,1,0,3,2],[2,3,0,1,4,5]]», Finite.of_fintype _, by decideFin!⟩
@@ 4.1. Problems of scale in mathematical collaboration
In order to understand the scaling issues that can arise in large scale collaborations, it helps to revisit the mechanics of traditional mathematical collaborations and their limitations. While every collaboration is unique, there are some general patterns. A small number of contributors, usually under ten, who may know each other, join forces to tackle some class of problems. Typically the collaborators are almost all academics who share substantial amounts of common knowledge. They discuss the problem at hand together, typically with some shared written medium such as a whiteboard. After several rounds of discussion and refinement, different members of the collaboration come up with different pieces of a solution. These pieces are then put together via discussion and merging of write-ups over several iterations. Once the collaborators are reasonably confident about the correctness of their work, including theorem statements and proofs, they submit the paper for peer review. Thus the correctness of mathematical research relies on this basic cycle of discuss, solve, write, cross-check, and revise, followed by peer review. Ultimately the authors take responsibility for the contents of their research article. This joint responsibility for authorship is formally enshrined by mathematical societies. For instance, see point 4 of the EMS code of practice for joint responsibility [EMS_code_of_practice].
@@ 4.2. The Blueprint tool
The formalization of proofs is an act of careful engineering. It is therefore helpful to have a blueprint with detailed natural language lemmata, definitions, and proof sketches in Lean. In the Lean community it has been conventional to use the Lean blueprint tool by Patrick Massot et al. [leanblueprint]. The typical formalization project has a clearly defined set of target theorems, and the authors of the project work with a known proof, to produce a clear roadmap for the formalization. The Lean blueprint tool is capable of linking each piece of this natural language document to its Lean encoding, tracking the dependency of definitions and theorems, and progress through them, by producing a key coloured dependency graph. Thus the managers of the formalization project can not only organise the project to distribute tasks among contributors, but also track when various pieces of the formalization are complete.

### v2

@@ 1. Introduction
The purpose of this paper is to report on the Equational Theories Project (ETP) [^1] https://teorth.github.io/equational_theories/, a pilot project launched [65] in September 2024 to explore new ways to collaboratively work on mathematical research projects using machine assistance. The project goal, in the area of universal algebra, was selected [^2] The specific mathematical goal was inspired by the MathOverflow question “Is there an identity between the associative identity and the constant identity?”, posed on 17 July 2023. to be particularly amenable to crowdsourced and computer-assisted techniques, while still being of mathematical research interest.
@@ 1.1. Magmas and equational laws
An equational theory is a collection of equational laws; in this paper we will primarily be concerned with theories generated by a single such law, although it is certainly of interest to explore larger theories as well. Equational theories are one of the simplest non-trivial examples of a theory in the model-theoretic sense; they also arise in various areas of computer science, such as term rewriting systems [7], automated theorem proving [45], and in the Dolev–Yao model [23] of interactive cryptographic protocols.
@@
We also define a variant: we say that $\mathrm{E}$ entails $\mathrm{E}^{\prime}$} for finite magmas, and write ${\mathrm{E}}\models_{\mathrm{fin}}{\mathrm{E}^{\prime}}$, if every finite magma that satisfies $\mathrm{E}$, also satisfies $\mathrm{E}^{\prime}$}. Clearly, the relation ${\mathrm{E}}\models{\mathrm{E}^{\prime}}$ implies ${\mathrm{E}}\models_{\mathrm{fin}}{\mathrm{E}^{\prime}}$; but, as observed by Austin [6], the converse is not true in general.
@@
  • The central groupoid law $\mathrm{x}\simeq(\mathrm{y}\diamond\mathrm{x})\diamond(\mathrm{x}\diamond\mathrm{z})$ (E168) is an order-$3$ law introduced by Evans [26] and studied further by Knuth [35] and many further authors, being closely related to central digraphs (also known as unique path property digraphs), and leading in particular to the discovery of the Knuth-Bendix algorithm [36]; see [40] for a more recent survey.
@@
  • Tarski’s axiom $\mathrm{x}\simeq\mathrm{y}\diamond(\mathrm{z}\diamond(\mathrm{x}\diamond(\mathrm{y}\diamond\mathrm{z})))$ (E543) is an order-$4$ law that was shown by Tarski [66] to characterize the operation of subtraction in an abelian group; that is to say, a magma ${\mathcal{M}}=(M,\diamond)$ satisfies $\mathrm{E}543$ if and only if there is an abelian group structure on $\mathcal{M}$} for which $x\diamond y=x-y$ for all $x,y\in M$.
@@
  • In a similar vein, it was shown in [50] (see also [51]) that the order-$4$ law $\mathrm{x}\simeq(\mathrm{y}\diamond\mathrm{z})\diamond(\mathrm{y}\diamond(\mathrm{x}\diamond\mathrm{z}))$ (E1571) characterizes addition (or subtraction) in an abelian group of exponent $2$; it was shown in [47] that the order-$6$ law $\mathrm{x}\simeq(\mathrm{y}\diamond((\mathrm{x}\diamond\mathrm{y})\diamond\mathrm{y}))\diamond(\mathrm{x}\diamond(\mathrm{z}\diamond\mathrm{y}))$ (E345169) characterizes the Sheffer stroke in a boolean algebra, and it was shown in [29] that the order-$8$ law $\mathrm{x}\simeq\mathrm{y}\diamond((((\mathrm{y}\diamond\mathrm{y})\diamond\mathrm{x})\diamond\mathrm{z})\diamond(((\mathrm{y}\diamond\mathrm{y})\diamond\mathrm{y})\diamond\mathrm{z}))$ (E42323216) characterizes division in a (not necessarily abelian) group.
@@
Some further examples of laws characterizing well-known algebraic structures are listed in [45].
@@
The Birkhoff completeness theorem [7, Th. 3.5.14] implies that an implication ${\mathrm{E}}\models{\mathrm{E}^{\prime}}$ of equational laws holds if and only if the left-hand side of $\mathrm{E}^{\prime}$} can be transformed into the right-hand side by a finite number of substitution rewrites using the law $\mathrm{E}$. However, the problem of determining whether such an implication holds is undecidable in general [49]. Even when the order is small, some implications [^3] Another contemporaneous example of this phenomenon was the solution of the Robbins problem [44]. can require lengthy computer-assisted proofs; for instance, it was noted in [34] that the order-$4$ law $\mathrm{x}\simeq(\mathrm{y}\diamond\mathrm{x})\diamond((\mathrm{x}\diamond\mathrm{z})\diamond\mathrm{z})$ (E1689) was equivalent to the singleton law $\mathrm{x}\simeq\mathrm{y}$ (E2), but all known proofs were found with computer assistance. [^4] We improved such a proof to make it human-readable, see the blueprint of the ETP. Furthermore, for the finite magma implication relation ${\mathrm{E}}\models_{\mathrm{fin}}{\mathrm{E}^{\prime}}$, no analogue of the Birkhoff completeness theorem is available.
@@ 1.2. The Equational Theories Project
Such systematic determinations of implication graphs have been seen previously in the literature; for instance, in [59], the relations between $60$ identities of Bol–Moufang type were established, and in the blog post [70, §17], some initial steps towards generating this graph for the first hundred or so laws on our list were performed. However, to our knowledge, the ETP is the first project to study such implications at the scale of thousands of laws.
The ETP requires the determination of the truth or falsity of $4694$^{2}=$22\,033\,636$$ implications (for both arbitrary magmas and finite magmas), or $4694$\times($4694$-1)=$22\,028\,942$$ if the reflexive implications ${\mathrm{E}}\models\mathrm{E}$ are removed; while one can use properties such as the transitivity of entailment to reduce the work somewhat, this is clearly a task that requires significant automation. It was also a project highly amenable to crowdsourcing, in which different participants could work on developing different techniques, each of which could be used to fill out a different part of the implication graph. In this respect, the project could be compared with a Polymath project [28], which used online forums such as blogs and wikis to openly collaborate on a mathematical research problem. However, the Polymath model required human moderators to review and integrate the contributions of the participants, which clearly would not scale to the ETP which required the verification of over twenty million mathematical statements. Instead, the ETP was centered around a GitHub repository in which the formal mathematical contributions had to be entered in the proof assistant language Lean, where they could be automatically verified. In this respect, the ETP was more similar to the recently concluded Busy Beaver Challenge [18], which was a similarly crowdsourced project that computed the fifth Busy Beaver number $BB(5)$ to be 47 176 87047\,176\,870 through an analysis of about $180$ million Turing machines, with the halting analysis being verified in a variety of computer languages, with the final formal proof written in the proof assistant language Coq [68, 18]. One of the aims of the ETP was to explore potential workflows for such collaborative, formally verified mathematical research projects that could serve as a model for future projects of this nature.
@@ 1.3. Outcomes
In the course of completing the implication graph, some interesting new algebraic structures were discovered. One such example concerns the magmas satisfying $\mathrm{E}1485$, which we refer to as weak central groupoids as they contain the central groupoids (satisfying $\mathrm{E}168$) as a subclass. In [35] it was observed that all finite central groupoids have order equal to a perfect square $n^{2}$; empirically, we have found that finite weak central groupoids always have order $n^{2}$ or $2n^{2}$, although we have no rigorous proof of this claim; they also have a graph-theoretic interpretation analogous to the interpretation of central groupoids as digraphs with the unique path property. For these and other observations we refer the reader to the blueprint of the ETP.
@@ 1.4. Further directions
We refer the reader to the ETP blueprint for a more in depth discussion of several of the topics covered in this report, including a detailed treatment of particularly difficult equations and complex magma constructions, further treatment of term rewriting theory and equational law invariants, a list of “hard” implications that required particular effort to resolve in our project, and some implementation details of how tools such as automated theorem provers were utilized.
@@ 2. Notation and mathematical foundations
A theory is a collection $\Gamma$ of equational laws; we say that a magma $\mathcal{M}$} satisfies a theory, and write ${\mathcal{M}}\models\Gamma$, if every law in $\Gamma$ is satisfied by $\mathcal{M}$}. If $\mathrm{E}$ is an equational law, we write $\Gamma\models\mathrm{E}$ if every magma that satisfies $\Gamma$ also satisfies $\mathrm{E}$. A free magma ${\mathcal{M}}_{X,\Gamma}=(M_{X,\Gamma},\diamond)$ for such a theory $\Gamma$ and an alphabet $X$ is a magma satisfying $\Gamma$ together with a map $\iota_{X,\Gamma}\colon X\to M_{X,\Gamma}$ which is universal in the sense that every function $f\colon X\to{\mathcal{M}}$ to a magma $\mathcal{M}$} satisfying $\Gamma$ uniquely determines a homomorphism $\varphi_{f,\Gamma}\colon{\mathcal{M}}_{X,\Gamma}\to{\mathcal{M}}$ such that $\varphi_{f,\Gamma}\circ\iota_{X,\Gamma}=f$. This magma is unique up to isomorphism; a canonical way to construct it is as the quotient ${\mathcal{M}}_{X}/\sim_{\Gamma}$ of the free magma ${\mathcal{M}}_{X}$ by the equivalence relation $\sim_{\Gamma}$ given by declaring $w\sim_{\Gamma}w^{\prime}$ if $\Gamma\models w\simeq w^{\prime}$ [7, Theorem 3.5.6]. If $\Gamma=\{\mathrm{E}\}$ consists of a single law $\mathrm{E}$, we write ${\mathcal{M}}_{X,\mathrm{E}}$, $\sim_{\mathrm{E}}$, $\varphi_{f,\mathrm{E}}$ for ${\mathcal{M}}_{X,\{\mathrm{E}\}}$, $\sim_{\{\mathrm{E}\}}$, $\varphi_{f,\{\mathrm{E}\}}$ respectively.
@@ 3. Formal foundations
All proofs in the ETP were ultimately formalized in the proof assistant language Lean, though in many cases the proofs were first written in an informal human document, which was then incorporated into the human-readable blueprint [42] that accompanied the formalization. Many of the computer-assisted proofs were also first generated as computer output from a source other than Lean, such as an ATP, and later converted to a Lean proof by a separate program custom-written for this task.
@@
@[equational_result]
theorem _root_.Equation1437_not_implies_Equation4269 :
$\exists$ (G : Type) (_ : Magma G), Equation1437 G $\wedge$ $\lnot$ Equation4269 G := by
[…]
constructor
$\bullet$ intro x y z
simp [op, add_assoc]
$\bullet$ simp only [not_forall, op]
use (0, 0), (2, 0)
decide
@@
@[equational_result]
theorem «Facts from All4x4Tables [[1,2,3,4,5,0],[4,1,2,5,0,3],[3,0,5,2,1,4],
[0,5,4,3,2,1],[5,4,1,0,3,2],[2,3,0,1,4,5]]» :
$\exists$ (G : Type) (_ : Magma G) (_ : Finite G), Facts G [1316, 2863] [411, 680,
817, 1020, 1426, 2035, 2441, 2644, 2853, 2855, 2865, 2872, 2947, 3050,
3253, 3456, 4270, 4283, 4290, 4380, 4598, 4605, 4656] :=
[…]
@@ 4.1. Problems of scale in mathematical collaboration
In order to understand the scaling issues that can arise in large scale collaborations, it helps to revisit the mechanics of traditional mathematical collaborations and their limitations. While every collaboration is unique, there are some general patterns. A small number of contributors, usually under ten, who may know each other, join forces to tackle some class of problems. Typically the collaborators are almost all academics who share substantial amounts of common knowledge. They discuss the problem at hand together, typically with some shared written medium such as a whiteboard. After several rounds of discussion and refinement, different members of the collaboration come up with different pieces of a solution. These pieces are then put together via discussion and merging of write-ups over several iterations. Once the collaborators are reasonably confident about the correctness of their work, including theorem statements and proofs, they submit the paper for peer review. Thus the correctness of mathematical research relies on this basic cycle of discuss, solve, write, cross-check, and revise, followed by peer review. Ultimately the authors take responsibility for the contents of their research article. This joint responsibility for authorship is formally enshrined by mathematical societies. For instance, see point 4 of the EMS code of practice for joint responsibility [64].
@@ 4.2. The Blueprint tool
The formalization of proofs is an act of careful engineering. It is therefore helpful to have a blueprint with detailed natural language lemmata, definitions, and proof sketches in Lean. In the Lean community it has been conventional to use the Lean blueprint tool by Patrick Massot et al. [42]. The typical formalization project has a clearly defined set of target theorems, and the authors of the project work with a known proof, to produce a clear roadmap for the formalization. The Lean blueprint tool is capable of linking each piece of this natural language document to its Lean encoding, tracking the dependency of definitions and theorems, and progress through them, by producing a key coloured dependency graph. Thus the managers of the formalization project can not only organise the project to distribute tasks among contributors, but also track when various pieces of the formalization are complete.